OpenStudy (anonymous):

solve for; e^x = 1/3

OpenStudy (anonymous):

\[e^x=1/3\] \[e^x=e^{\log_{e}1/3}\] \[x=\log _{e}1/3\] Put that into a calculator

OpenStudy (anonymous):

ok, thnx :)